The magnetic field strength at the center of a circular coil with N turns is B. If the number of turns is double but the current and radius are each half, the magnetic field strength at the center would be 


1
Expert's answer
2022-02-25T11:12:00-0500

Answer

Magnetic field at center of coil is given

By

"B=\frac{N\mu_0 I}{2R}"

If

number of turns is double but the current and radius are each half

Then

magnetic field strength at the center would be

"B'=\frac{2*2N\mu_0 I}{2R}\\=4B"

Magnetic field become four times as earlier.